OpenStudy (anonymous):

Find the equivalent equation in rectangular form. r(cos(theta)-sin(theta))=4

OpenStudy (amistre64):

z^1 ?

OpenStudy (amistre64):

cos(t)-i sin(t) = 4 maybe

OpenStudy (amistre64):

what is rectangular form?

OpenStudy (amistre64):

cos(t) = x sin(t) = y r tends to be a radial measure of this is going from polar to a more normal equation

OpenStudy (amistre64):

r = sqrt(x^2+y^2)

OpenStudy (amistre64):

\[\sqrt{x^2+y^2}(x-y)=4\] should be it then

OpenStudy (asnaseer):

\[x=r\cos(\theta)\]\[y=r\sin(\theta)\]so:\[r(\cos(\theta)-\sin(\theta))=4\]becomes:\[x-y=4\]

OpenStudy (amistre64):

..... soo close :)

OpenStudy (asnaseer):

:)